A block of material with a cross-sectional area of 15 cm2 and length of 8.0 cm is at hand. A temperature difference of 30 ̊C is established and maintained across the block. Find the rate of heat flow through it if: (2 marks)

  1. The material is copper (K = 397 Js-1m-1C-1).

  2. The material is air (K = 0.0234 Js-1m-1C-1).

  3. The material is wood (K = 0.08 Js-1m-1C-1).
Answer #1

Given,

A = 15 cm^2 = 0.0015 m^2 ; t = 8 cm = 0.08 m ; T = 30 deg C

We know that

dQ/dt = k A dT/t

R1 = 397 x 0.0015 x 30/0.08 = 223.3125 J/s

Hence, R1 = 223.3125 J/s

R2 = 0.0234 x 0.0015 x 30/0.08 = 0.0132 J/s

Hence, R2 = 0.0132 J/s

R3 = 0.08 x 0.0015 x 30/0.08 = 0.045 J/s

Hence, R3 = 0.045 J/s